UT 20131009UT FLAO+PISCES DX

Observer: Guerras,JC,Christou(LBTO),Arcidiacono,C, Boutsia, K, (INAF)
Telescope Operator:Bechetti G
SW Support:
AO Support: JChristou
Telescope Support: JLittle
Instrument Support: Edwards,M

Summary

Problems at the beginning of the night with the FLAO. Z-stage was not moving. JC and MWagner had to go up and open up the system. 2.0hr lost.

Project_RUN Objects Filter ExpTime Notes
Beatty kELT-1 H 1600sec COMPLETED - a set with FeII (0.8s) because some of the H (0.8s) images were saturated
Guerra NGC404 H/K/J 600s/360s/840s COMPLETED - star for PSF and photometric calibration
Cesaroni IRAS05173 Ks 640s 80% of this band - the field was difficult to recognize - loop was unstable because we're facing the wind
Beatty KC04C02 H 120sec 50% - we had to abort because of twilight - the target is saturated - ask instructions before repeating

Details

02:25 The TO is running pointing and collimation

02:29 We slew to the Cusano-Corot targets - elev. 48
  • a lot of astigmatism during first collimation

02:40 send the first preset to Cusano_652345526
  • preset, CUS_652345526, MODE=ACE, GS=2, TS=0, RS=0, PA=30
  • first preset - Z-stage stuck!!
  • we changed PA in the script and we send the preset again - still stuck
  • Carmelo changes manually the stages positions to see if it works - it doesNOT!
  • the wind is blowing straight on the shell (gusts up to 17m/s) putting additional strain...
  • we give up with this target for now

02:55 We slew to KELT_1
  • preset, KELT_1, MODE=ACE, GS=9, TS=0, RS=0, PA=225.0
  • the same problem also on this target - Z stage stuck - Carmelo il manually moving the other stages trying to free it ...

02:59 Trying to unblock the Z-stage we move the telescope zenith - nothing happens
  • TO is closing the chambers - moving telescope to horizon
  • Carmelo and JC are going to the gallery for a hard reset of the FLAO (manually push the motors)

03:33 They tried to move it changing the rotator angles and manually moving the stages from the gui...
  • at one angle Pisces was upside down and it started spilling
  • this did not work - Elliot is filling up pisces
  • FLAO has been turned off and they will remove the cover

03:50 The Zstage seems to be moving - JC is putting back the cover
  • Elliot also had pisces filled up in the meantime!

04:00 The TO is opening the dome

04:12 Preset to KELT-1 - now the guide probe is stuck - the TO needs to power cycle the GCS!!

04:22 Preset to KELT-1 again
  • preset, KELT_1, MODE=ACE, GS=18, TS=0, RS=0, PA=180.000
  • the shell ripped
  • repeat preset - trying to close the AO loop

04:31 preset failed - problems with the FLAO - JC is power-cycling the AO

04:45 the AOS is up again - repeat KELT-1 preset
  • CCD47 was still not ready
===> Error in irc_presetTelescope err = Resmsg[0] = PresetTelescope? result status: Error err = Resmsg[1] = AOSR err = Resmsg[2] = Stop[46] complete err = Resmsg[3] = AOSR err = Resmsg[4] = Timeout waiting for techviewctrl.R.STATUS (-5001) TIMEOUT_ERROR err = Resmsg[5] = PresetAO? [47] failed - Timeout waiting for techviewctrl.R.STATUS (-5001) TIMEOUT_ERROR

04:47 send preset again - this time it worked!
  • preset, KELT_1, MODE=ACE, GS=18, TS=0, RS=0, PA=180.000

05:02 after centering the source we start with the dither patterns - short

05:16 short dither pattern with FeII because with 0.8s some of the H frames were saturated

05:25 long dither pattern - filter H - exptime 10s

seeing 1.1"

05:57 repeat dither pattern with better offsets in order to minimize the corquad effect.
  • 198 frames of 10sec obtains

06:37 Slew to NGC404 - loop closed with bin4

06:50 dither pattern with exptime 60s - H filter

07:07 send preset again - because we had to change the PA to keep pisces from spilling
  • failed to close the AO loop

07:17 send preset again - it was difficult to close the loop - we had to send the preset several times because of failures

07:33 start dither pattern again in H filter with larger offsets for the sky
  • the previous sky was not flat enough, since there seemed to be still inside the galaxy halo.
  • with the large offset (900") we have a lot of warnings but the script did not fail

07:49 start dither pattern in filter K

08:10 start dither pattern in filter J

08:49 slew at PSF_N404 for PSF and photometric calibration
  • obtain images in J,K,H

09:39 The TO is adjusting the pointing before we slew to the next target

09:46 send preset for Cesaroni - IRAS05137
  • preset failed - no star found
  • adjust collimation and send preset again
  • offsetting the target so that the jets are inside the FOV
  • it takes some time to recognize the field because we do not know the pisces orientation and the jets are extended and the correct angle has to be set in order to have them in the FOV

11:10 dither pattern in K band on IRAS05137

11:25 the AO loop opened - send the preset again

11:31 right after the preset the loop opened

11:34 send preset again

11:40 started again the dither in the K band
  • shell ripped - we are facing the wind and the loop is unstable

11:47 start again dither pattern in K

10:52 abort the script - there is something weird wit the loop that needs to be checked

10:54 started the dither again - with 3 frames at each position, just to see if it makes a difference
  • probably taking one image per position is not a very good strategy, because the system doesn't have enough time to stabilize...

12:00 the AO loop is open - we need to send again the preset

12:15 we aborted the preset because pisces was spilling
  • in the meantime we take sky frames
  • 63/115 useful science images + 36 useful sky images

12:15 we slew to KC04C02 - the rotator is stuck

12:27 preset to KC04C02

12:54 abort because the loop opened

12:57 moving to zenith for flats
  • flats in the H band low/high

13:23 Close dome - darks: 0.8s, 10s, 120s, 60s, 80s

Vibration detected @30Hz

It has been detected events of vibrations@30H during the night. Attach a sample of the events.

CommandLineCutAndPaste

-- KonstantinaBoutsia - 09 Oct 2013
  • Vibration@30Hz:
    vibration_053224-053314_2.png
Topic revision: r10 - 13 Oct 2013, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback